# Exploring the Potential of Large Language Models for Improving Digital Forensic Investigation Efficiency

### v2

@@ 5.1 Challenges for LLMs in Digital Forensics
To optimise results, LLMs likely need to be trained with specific forensic data (i.e., previous case data) to achieve the best results. Given the complexity and variation of the cases, it is questionable how good the training data are and whether there are sufficient data [158]. Any bias in the training data can lead to skewed interpretations and unjust outcomes.
Investigative findings and conclusions are also highly based on investigator experience and intuition, which are difficult to replicate using LLMs. Initially, they will only be effective for certain subtasks within digital forensics, raising questions about their superiority over deterministic programs in certain contexts. For instance, data may be parsed/converted using regular software or an LLM. While the former works deterministically, the latter is more general, but requires the practitioners to validate the output. This leads to the fundamental necessity for explainability (i.e., not using blackboxes) for investigations. The explainability of LLM-generated results remains a daunting challenge, as understanding the reasons for their, often non-deterministic, results is often elusive [157].
@@
Maintaining the energy and resource intensive infrastructure required to train and deploy LLMs is a significant financial burden, making it impractical for smaller forensic laboratories to utilise such technologies. Centralised systems may be an option, but require clear guidelines on how data can/may be shared and exchanged. Although LLMs can serve as valuable tools to support forensic investigations, it should be recognised that they currently function best as an aid, not a substitute for human expertise [159]. There is a risk that people may place too much trust in the results generated by LLMs (over-reliance), which could lead to complacency and overlook the need for detailed human expert analysis and validation.
@@
To mitigate the potential misuse of LLMs, many LLMs undergo censorship [160, 58]. While this censorship may serve as a preventive measure against unethical use, it can pose challenges in the field of DF. For instance, if an investigator seeks evidence related to ‘drugs’ or evidence of other illegal material, censorship of the LLM may restrict access to accurate information pertaining to the investigator’s query. This limitation underscores the need for a nuanced approach to censorship in LLMs, balancing ethical considerations with the imperative of facilitating effective forensic investigations. In addition, the censorship of generic, publicly accessible LLMs further supports the argument for a discipline specific DF LLM.
@@
Finally, ethical and legal considerations must also be discussed. Determining accountability in cases where LLMs produce false information or are compromised by hacking. Clarifying responsibilities between developers, users, and regulators is crucial to establish a framework for accountability. If LLM generated digital forensics (DF) results lead to incorrect information, the responsibility may lie with both the developers, for ensuring the model’s accuracy, and the users, for appropriately interpreting and validating the results.
@@ 5.2 Risks of Integration
The integration of LLMs within the DF process comes with inherent risks – extra to those general LLM limitations outlined in Section 3.7. Notably, in the examination, analysis, and reporting phases, the utilisation of LLMs introduces the risk of producing inaccurate information, primarily due to the phenomenon of inheritance hallucinations associated with these models [157, 159]. Additionally, the biases and obscurities present in an inheritance model can significantly impact the performance of a DF-focused LLM – potentially leading to the unacceptable generation of biased or inaccurate information within the DF process.
@@
Mitigating these changes and risks can be challenging, particularly in scenarios involving adopting country-specific legal systems. However, there are potential strategies to address technical challenges such as hallucination, censorship, and substantial infrastructure costs.
One solution to mitigate hallucination was suggested by Ji et al. [162], who proposed an interactive self-reflection method for generated knowledge and answers – an approach that has been demonstrated promise. Another hallucination mitigation method is using RAG, which provides a larger knowledge base to the LLMs to minimise unknown information [132]. Other methods such as knowledge graphs, bias detection mechanisms, active learning methods for LLMs, supervised fine-tuning strategies, hallucination mitigation frameworks and new decoding strategies can also help mitigate hallucination to some extent [163, 164].
@@
The high infrastructure costs associated with these models can be mitigated by employing Data Forensics as a Service (DFaaS) platforms such as Hansken. With DFaaS, investigators need only input queries related to their investigations using personal computers, while the platform manages the model maintenance and computational demands [29].
@@ 6 Conclusion
The convergence of LLMs with an array of technologies represents an exciting synergy. Although the utilisation of LLMs in the realm of DF is still in its nascent stages, there is evidence of their substantial potential to significantly augment the efficiency of investigations. The exploration of investments for LLMs across the entire DF process is considered, with the aim of improving the productivity and efficiency of investigations. Additionally, the integration of LLMs into current DF tools is posited to reduce user training times, as these models comprehend natural language input and provide output accordingly. In the dynamic landscape of LLM applications for DF, promising avenues for further exploration and advancement unfold.
@@
While the surge in LLM research is promising, it is crucial to balance enthusiasm with an awareness of existing challenges. The propensity for LLMs to produce hallucinations highlights the need for human oversight in critical decision-making processes, underscoring the irreplaceable value of human judgement, intuition, and expertise. A notable limitation is the language dependency issue, as most LLMs are predominantly trained on English data, reducing their effectiveness with non-English content. Additionally, the deployment of LLMs in DF involves significant costs related to infrastructure for processing evidence. Questions also arise about the validation of task correctness and quality when automated by LLMs, as well as the legal and professional acceptance of results obtained with limited human intervention.
@@
The trustworthiness of LLMs remains a debatable issue that requires careful attention. It is crucial to establish clear boundaries and measures to define LLM trustworthiness. Addressing this will be a key aspect in the field of DF, ensuring that LLMs can be relied upon for accurate and secure analysis, with explainability of their operations being paramount.
@@
Integrating LLMs with automated agents offers a promising path to automating DF processes, potentially allowing multiple cases to be handled concurrently for more timely and precise outcomes. This integration could significantly streamline investigations. Future research should explore the role of LLMs and AI in DF decision-making. It is essential to focus on validating LLM generated outputs to ensure their scope, accuracy, reliability, and trustworthiness in investigations. Further studies comparing DF outcomes with and without LLM integration are critical, as they could highlight LLMs’ benefits and controlled applicability in DF and similar fields.
@@
In essence, while LLMs offer exciting prospects for the future of digital forensics, a balanced approach that integrates their strengths with human oversight is essential for harnessing their full potential. Inevitably, LLM facilitated DF processes will themselves become the focus of future investigation.

### v3

@@ 5.1 Challenges for LLMs in Digital Forensics
To optimise the results, the LLMs will likely need to be trained with specific forensic data (i.e., previous case data) to achieve the best results. Given the complexity and variation of the cases, it is questionable how good the training data are and whether there are sufficient data [165]. Any bias in training data can lead to skewed interpretations and unjust outcomes. This problem of bias can be mitigated by using diverse and representative datasets during the training phase, e.g., datasets that come from diverse sources, different case types and geographic regions. Furthermore, techniques such as data filtering, distribution reconstruction, rebalancing, regularisation, and prompting can be implemented to actively identify and correct biased patterns in the base data sets of the model and its outputs [166, 167]. These techniques involve adjusting model weights or incorporating fairness constraints during training to reduce the likelihood of biased predictions. Regular audit of the model’s performance against fairness benchmarks is also crucial to ensure that it remains unbiased over time [168].
The experience level of investigators and the practical strategies employed in conducting investigations are challenging to replicate with LLMs. Initially, LLMs can excel in assisting with certain subtasks, such as parsing and data conversion, tasks in which output can be easily verified. However, when it comes to more interpretative or inferential tasks, LLMs’ lack of inherent transparency introduces explainability challenges. Unlike deterministic software, whose logic can be easily traced, LLMs often act as black boxes, making it difficult to validate and understand the rationale behind their conclusions, particularly when these outputs extend beyond straightforward parsing into areas requiring judgement and reasoning. This underscores the importance of explainability in the application of LLMs to forensics, where understanding the basis of each result is crucial for accuracy and accountability [164].
Publicly hosted and maintained LLMs are generally unsuitable for casework due to the sensitivity of the evidence and information involved, which require strict privacy and security controls that cannot be reliably ensured on public platforms. Furthermore, managing the substantial infrastructure needed for LLM training and deployment is both energy and resource intensive, presenting a financial hurdle, especially for smaller forensic laboratories with limited budgets. Although methods like retrieval-augmented generation (RAG) or prompt engineering can reduce some of the computational load by tailoring responses with existing models, they still require powerful GPU resources to effectively run these models, adding to the cost and accessibility barriers. Centralised systems could be a viable option, but they require well-defined guidelines for data sharing and stringent security standards to safeguard sensitive information.
@@
Although LLMs can serve as valuable tools to support forensic investigations, it should be recognised that they currently function best as an aid, not a substitute for human expertise [169]. There is a risk that people may place too much trust in the results generated by LLMs (over-reliance), which could lead to complacency and overlook the need for detailed human expert analysis and validation.
@@
To mitigate the potential misuse of LLMs, many LLMs are subjected to censorship [170, 171]. Although this censorship may serve as a preventive measure against unethical use, it can pose challenges in the field of DF. For example, if an investigator seeks evidence related to ‘drugs’ or evidence of other illegal material, censorship of the LLM may restrict access to accurate information related to the investigator’s query. This limitation underscores the need for a nuanced approach to censorship in LLMs, balancing ethical considerations with the imperative of facilitating effective forensic investigations. In addition, the censorship of generic, publicly accessible LLMs further supports the argument for a discipline-specific DF LLM.
@@
Finally, ethical and legal considerations must also be discussed. Determining accountability in cases where LLMs produce false information or are compromised by hacking. Clarifying responsibilities between developers, users, and regulators is crucial to establish a framework for accountability. If LLM generated DF results lead to incorrect information, the responsibility may lie with both the developers, for ensuring the model’s accuracy, and the users, for appropriately interpreting and validating the results.
@@ 5.2 Risks of Integration
The integration of LLMs within the DF process comes with inherent risks, in addition to the general LLM limitations outlined in Section 3.7. In particular, in the examination, analysis, and reporting phases, the use of LLMs introduces the risk of producing inaccurate information, primarily due to the phenomenon of inheritance hallucinations associated with these models [164, 169]. Additionally, the biases and obscurities present in an inheritance model can significantly impact the performance of a DF-focused LLM – potentially leading to the unacceptable generation of biased or inaccurate information within the DF process.
Hallucinations in LLMs present a considerable risk, as they can produce information that appears credible but is incorrect. This can lead law enforcement authorities to form invalid assumptions and make flawed decisions based on unreliable results. Additionally, inherent biases in LLMs can influence investigative outcomes, which could affect the fairness and integrity of legal procedures. Data privacy concerns are also prominent, as sensitive information confidentiality may be compromised when using LLMs in DF processes. Together, these factors present substantial challenges to the reliable and ethical application of LLMs.
@@
Mitigating these changes and risks can be challenging, particularly in scenarios that involve adopting country-specific legal systems. However, there are potential strategies to address technical challenges such as hallucinations, censorship, and substantial infrastructure costs.
One solution to mitigate hallucination was suggested by Ji et al. [173], who proposed an interactive self-reflection method for generated knowledge and answers, an approach that has shown promise. Another method of reducing hallucinations is the use of RAG, which provides a larger knowledge base for LLMs to minimise unknown information [135]. Other methods such as knowledge graphs, bias detection mechanisms, active learning methods for LLMs, supervised fine-tuning strategies, hallucination mitigation frameworks, and new decoding strategies can also help mitigate hallucinations to some extent [174, 175].
@@
The high infrastructure costs associated with these models can be mitigated by employing Data Forensics as a Service (DFaaS) platforms such as Hansken. With DFaaS, investigators only need to input queries related to their investigations using personal computers, while the platform manages the model maintenance and computational demands [20].
Despite these promising integration risks, the use of LLMs may face limitations in adaptability. The performance of an LLM is inherently tied to the dataset on which it was trained, which means that its ability to respond to new or emerging information is constrained. For example, if an LLM is tasked with identifying possible malware in a system, it may struggle to detect newer malware variants that were not part of its training data [31]. To mitigate such issues, LLMs need to be fine-tuned frequently, which poses its own challenges due to the significant computational power required for such operations.
@@ 6 Conclusion
The convergence of LLMs with an array of technologies represents exciting synergy. Although the utilisation of LLMs in the realm of DF is still in its nascent stages, there is evidence of their substantial potential to significantly increase the efficiency of investigations. The exploration of investments for LLMs across the entire DF process is considered, with the aim of improving the productivity and efficiency of investigations. Additionally, the integration of LLMs into current DF tools is posited to reduce user training times, as these models comprehend natural language input and provide output accordingly. In the dynamic landscape of LLM applications for DF, promising avenues for further exploration and advancement unfold.
@@
Although the surge in LLM research is promising, it is crucial to balance enthusiasm with awareness of existing challenges. The propensity of LLMs to produce hallucinations highlights the need for human oversight in critical decision-making processes, underscoring the irreplaceable value of human judgement, intuition, and expertise. A notable limitation is the language dependency issue, as most LLMs are predominantly trained on English data, reducing their effectiveness with non-English content. Furthermore, the deployment of LLMs in DF involves significant costs related to the infrastructure to process evidence. Questions also arise about the validation of task correctness and quality when automated by LLMs, as well as the legal and professional acceptance of results obtained with limited human intervention.
@@
The trustworthiness of LLMs remains a debatable issue that requires careful attention. It is crucial to establish clear boundaries and measures to define LLM trustworthiness. Addressing this will be a key aspect in the field of DF, ensuring that LLMs can be trusted for accurate and secure analysis, with the explainability of their operations being paramount.
@@
Integrating LLMs with automated agents offers a promising path to automating DF processes, potentially allowing multiple cases to be handled concurrently for more timely and precise outcomes. This integration could significantly streamline investigations. Future research should explore the role of LLMs and AI in the decision making of DF. It is essential to focus on validating LLM generated outputs to ensure their scope, accuracy, reliability, and trustworthiness in investigations. More studies comparing DF outcomes with and without LLM integration are critical, as they could highlight the benefits of LLMs and the controlled applicability of LLMs in DF and similar fields.
A future use case involves developing forensic-specific LLMs fine-tuned for automated examinations. These models could be optimised for script generation to support investigations where no existing tools are available, allowing forensic analysts to create customised solutions on demand. Integrating AI agents with these models could streamline evidence handling by allowing investigators to perform complex queries more intuitively, such as retrieving all messages from a specific date without the need to craft regular expressions.
@@
In essence, while LLMs offer exciting prospects for the future of digital forensics, a balanced approach that integrates their strengths with human oversight is essential for harnessing their full potential. Inevitably, LLM-facilitated DF processes themselves will become the focus of future investigation.